If you travel 320miles in 1.6hrs calculate speed in: milometers hr, feet per second, kilometer per hour, centimeters per second
1
Expert's answer
2012-08-21T09:45:44-0400
V = 320/1.6 miles per hour
= 200 miles per hour
= 1.609344*200 kilometers per hour
= 321.8688 kilometers per hour
= 321.8688*100000/3600 centimeters per second
= 8940.8 centimeters per second
= 8940.8/30.48 feet per second
= 293.(3) feet per second.